# I Introduction

### v2

@@ I Introduction
A powerful indirect probe to constrain or measure weak magnetic fields makes use of $\gamma$-ray observations of extragalactic sources (see, e.g., Ref. [6] for a review). Very-high energy (VHE, $E>100$ GeV) $\gamma$-ray photons can be absorbed during their propagation by the extragalactic background light (EBL) via the $\gamma$-$\gamma$ pair production process ($\gamma$$\gamma\rightarrow e^{+}$$e^{-}$) [7, 8]. The created pairs inverse Compton (IC) scatter photons from the cosmic microwave background (CMB) up to $\gamma$-ray energies, initiating an electromagnetic cascade. If an IGMF is present, the pairs are deflected, which induces several observational signatures of the cascade emission, such as an additional spectral component and a $\gamma$-ray halo around otherwise point-like sources. The absence of these features in $\gamma$-ray observations of active galactic nuclei (AGNs) with their jets closely aligned to the line of sight (so-called blazars) has led to lower limits on the IGMF in the range of 10-15–10-17 G [e.g., 9, 10, 11, 12, 13, 14]. The results depend on various assumptions such as the intrinsic VHE spectral shape, the AGN duty cycle, as well as the coherence length of the IGMF. Recently, a lower bound of the order of $\sim$ 10-17 G on IGMF strength was derived, which took into account the variability pattern of the source in the VHE band [15].
The discovery of the VHE emission from gamma-ray bursts (GRBs) [16, 17] offers a unique opportunity to constrain the IGMF. A non-zero IGMF would cause a time delay of the cascade emissions, producing observable $\gamma$-ray emission even after the GRB afterglow has ceased [18, 19, 20, 21, 22]. This approach has the advantage that it does not require assumptions on the source’s duty cycle or variability pattern as these quantities are directly determined for GRBs. Several attempts have been made to measure the IGMF using this pair-echo emission using the first GRB detected in the VHE band, GRB 190114C [23, 24, 25, 26]. The authors of Ref. [23] excluded IGMF strengths lower than $\sim$ $10^{-19.5}$ G for a coherence length $\leq$ 1 Mpc, which was however, not confirmed in other studies [24, 25]. Instead, no limits could be derived. These discrepancies can be explained with the different assumptions made for the intrinsic VHE spectral shape of the GRB and shifting from semi-analytical to cosmic-ray propagation codes for the calculation of the cascade spectral energy distributions (SEDs). Therefore no IGMF constraint can be inferred from this GRB.
@@ II The VHE intrinsic spectrum of GRB 221009A
The GRB 221009A was an extremely energetic ($E_{\gamma}^{\rm iso}\sim 10^{55}$ erg) event detected with the Fermi gamma-ray burst monitor (GBM) on 2022 October 9 at $T_{0}=13:16:59.99$ UT [33]. With a redshift $z=0.151$, corresponding to a luminosity distance $D_{L}\sim 720$ Mpc, its isotropic luminosity is estimated to be $L_{\gamma}^{\rm iso}\simeq 10^{53}$ erg s-1 [34]. The event has been subsequently detected with the Fermi LAT in the GeV domain [35], and by the Swift-BAT and XRT instruments in X-rays [36]. Observations started 53 min and 55 min after $T_{0}$, respectively.
At the trigger time $T_{0}$, the event was in the field of view of LHAASO, which detected more than $6.4\times 10^{4}$ photons within the first 3000 seconds at $E_{\gamma}>200$ GeV [37]. LHAASO is a hybrid array of VHE detectors, composed of both water Cherenkov photo-multipliers (WCDA) and solid-state scintillators (KM2A) that are sensitive to $\gamma$ rays above a few TeV [38]. The intrinsic differential spectra in the range 200 GeV-7 TeV, derived within 5 temporal bins ranging from $T_{0}+5$ s to $T_{0}+1774$ s, do not show any evidence of energy cutoff up to 5 TeV [30]. The spectra are compatible with power laws (PLs) with indexes in the range 2.1–2.5. The KM2A observations reveal $\gamma$ rays with energies up to 13 TeV [31].
@@
Figure 2: Left: Simulated SEDs of the pair-echo emission expected between 3.3 and 365 days after $T_{0}$ with the same injection spectrum as in Fig. 1 compared with the Fermi-LAT upper limits. Right: Likelihood profiles sampling the magnetic field for different time intervals. The astrophysical GRB afterglow model assumes an ad-hoc power-law spectrum with $\Gamma=2$. Solid lines represent cubic spline interpolations between the derived likelihood values. The results were derived using the same set of simulations as in Fig. 1.
@@ III Simulation of the pair-echo emission
To produce the pair-echo SEDs in different time windows we employ the Monte-Carlo code CRPropa 3 (version 3.2), which simulates the development of the pair-echo emission given a primary VHE photon spectrum injected in the intergalactic medium [32]. The source is located at the center of a sphere with radius $D$ equal to the comoving distance to the source. We assume that the source emits VHE photons in the range 0.1-30 TeV according to Eq. (1), and re-weight the SED accordingly when a cutoff is adopted. These photons are injected considering an aperture cone of 1.6, as inferred from the break in the LHAASO light curve [30]. The choice of the cone aperture is based on the GRB emission model adopted by the LHAASO collaboration. We note that a larger opening angle could affect the cascade flux only for large time delays. However, for the time delays considered in this work, this parameter does not significantly impact the results.
Examples of the derived pair-echo light curves and SEDs are shown in Fig. 1 and Fig. 2. In Fig. 1, we show the cascade flux for $E>100$ MeV as a function of time for different IGMF strengths for a cutoff in the intrinsic spectrum of 7 TeV. The left panel of Fig. 2 shows the SED averaged between 3.3 days and 365 days. The lower value of 3.3 days corresponds to the last detection from the GRB as reported by Fermi LAT [39]. As expected, both figures show that the cascade flux drops with increasing magnetic field, since the time delay of the cascade photons increases with increasing IGMF strength. Furthermore, the peak of the cascade flux moves towards higher energies with increasing IGMF. This is due to the longer time delay of low energy photons as low energy pairs have a larger Larmor radius for higher IGMF values. We provide further details on the CRPropa simulations in Supplemental Material [40].
IV Fermi-LAT limits on the pair echo-emission
The Fermi LAT is a pair-conversion telescope that detects $\gamma$ rays from 20 MeV to beyond 300 GeV [41]. We select events between 100 MeV and 1 TeV, within a $10^{\circ}\times 10^{\circ}$ region centered on GRB 221009A, with a zenith angle cut at $90^{\circ}$. We divide the dataset in several time bins following the main analysis of the Fermi-LAT Collaboration [39], expanding it up to 1 year after $T_{0}$ [40].
The likelihood of observing the data $\mathcal{D}_{ij}$ in the $i$-th time bin and $j$-th energy bin follows a Poisson likelihood $\mathcal{L}(B,\bm{\theta}_{i}|\mathcal{D}_{ij})$. This assumes an IGMF strength $B$ and spectral parameters $\bm{\theta}$ for both the background sources and potential GRB afterglow emission. Background sources can be nearby point sources or diffuse backgrounds. The log-likelihood ratio test can then be written as
@@ IV Fermi-LAT limits on the pair echo-emission
where $\hat{\hat{\bm{\theta}}}_{i}$ denotes the spectral parameters maximizing $\mathcal{L}_{i}$ for a fixed IGMF field strength in the $i$-th time bin, whereas $\hat{\bm{\theta}}_{i}$ and $\hat{B}$ are the parameters that maximize $\mathcal{L}_{i}$ unconditionally. Note that, for time bins earlier than $T_{0}+3.3\,$ days, we additionally included a point source with a power-law spectrum with index $\Gamma=2$ and free normalization to account for the GRB afterglow emission [40].
We show the results of the log-likelihood ratio test in the right panel of Fig. 2. No IMGF strength is preferred. Instead, we are able to rule out IGMF strengths at 95 % confidence for which $\lambda>2.71$. Combining all time and energy bins, this excludes $B<2.5\times 10^{-17}$ G. The most stringent constraint comes from the latest time bin (up to $T_{0}+1$ yr), which rules out $B<2.2\times 10^{-17}$ G. Constraints from the individual time bins are reported in [40].
@@ V Discussion and conclusions
Our results depend on the chosen high-energy cutoff of the injection spectrum of VHE $\gamma$ rays as well as the assumed GRB afterglow emission in the early time bins. Increasing the cutoff energy to 13 TeV, corresponding to the highest energy reported with the LHAASO KM2A detectors rules out $B<3.5\times 10^{-17}$ G. This is expected, as photons injected at even higher energies will increase the amount of pair-echo photons in the Fermi-LAT energy band. If no cut-off is applied to the log-parabola spectrum the constraints further improve and rule out an IGMF with $B<5\times 10^{-17}$ G. These results assume that an afterglow component can account for the GeV signal, fitting a power-law with $\Gamma=2$ to the flux observed by the LAT in each time bin. To estimate the impact of assuming such an ad-hoc component, we alternatively test a scenario where we replace the simple power-law model with a physically motivated SSC model. This model is set to reproduce the LHAASO data, and only employed to predict the GeV flux in the $i$-th time bin [40]. The limits weaken marginally, and an IGMF with $B<1.4\times 10^{-17}\,$G is ruled out for a high-energy cutoff at 7 TeV. Finally, we also considered the scenario of no astrophysical afterglow. In this case, a preference for the cascade is found in the early time bins in order to explain the Fermi LAT detection. However, different time bins prefer different values of the IGMF, which are all incompatible with the constraints from the latest time bin extending up to $T_{0}+1\,$yr. Therefore, we do not regard this as evidence for an IGMF.
Figure 3: The IGMF parameter space ($B$, $\ell_{B}$). Theoretically preferred regions by electro-weak (EWPT) or QCD phase transitions (QCDPT), or by Galactic winds and induced by cosmic-ray (CR) streaming are shown in blue [3]. Previous constraints using LAT observations of GRB 221009A, as well as constraints from $\gamma$-ray observations of blazars are shown as gray lines and gray shaded regions with arrows. Our constraints at the 95 % confidence level are shown in orange.
In principle, the choice of the EBL model could affect the amount of cascade flux. In this work, we used the EBL model by [42]. However, as we show in the Supplemental Material [40], the cascade flux level does not vary significantly across different EBL models, so the IGMF limit is not sensitive to the particular choice of model.
@@
Our results are significantly more constraining (by more than an order of magnitude) than previous analyses also using Fermi-LAT observations of GRB 221009A [27, 28, 29], see Fig. 3. The most significant factor concerns the time window considered for searching the pair echo. Previously, the cascade signal was searched for up to 20 days [29], 90 days [28] or 8 months [27] after $T_{0}$. Furthermore, in our analysis we make full use of both spectral and temporal information given the chosen time binning.
Given the exceptionally high flux in the VHE band, it is plausible that Fermi-LAT observations could still be sensitive to the pair-echo emission at later times of more than one year. We assess this possibility by extending our simulations and extrapolating the LAT exposure up to 5 years. Given the energy-dependent dilution in time of the signal, we find that the average flux expected within the 3-15 GeV band will tighten the constraints if several years of exposure are added, despite not reaching an additional order of magnitude [40].
The lower bound on the IGMF strength derived in this work holds as long as the correlation length $\ell_{B}$ of the IGMF is larger than the cooling distance of the electron-positron pairs for IC scattering, $l_{IC}\simeq 0.7\left(E_{\gamma}/~\rm{TeV}\right)^{-1}$ for primary $\gamma$ rays with energy $E_{\gamma}$. Considering $E_{\gamma}\sim 10\,$TeV, we find $\ell_{B}<l_{\mathrm{IC}}$ for $\ell_{B}\sim 0.1\,$Mpc, under the assumption that the electron and positron have both half the energy of the $\gamma$ ray. For lower values of the coherence length, the lower bound on the IGMF strength should scale as $\ell_{B}^{-1/2}$ for a Kolmogorov turbulence spectrum [44], as electron-positron stochastic deflections can be described as diffusion in angle (see Ref. [45] for the scaling with different turbulence spectra). Our result represents the best constraint so far on the IGMF employing the pair-echo technique, comparable to those obtained by studying blazars [15]. This opens the door for IGMF studies by means of TeV-detected GRBs to be as competitive as those done on persistent extragalactic VHE sources, while not suffering from assumptions on the duty cycles [14]. Moreover, our results are not affected by the possible development of plasma instabilities in the electron-positron beam in voids [46]. It has been argued that such instabilities could lead to an energy loss of the pairs on time scales shorter than the IC cooling time. However, the time required for such instabilities to grow is one order of magnitude larger than the time activity of the GRB 221009A, even with its large TeV luminosity [46, 30].

### v3

# Constraints on the intergalactic magnetic field from Fermi-LAT observations of GRB 221009A

@@ I Introduction
A powerful indirect probe to constrain or measure weak magnetic fields makes use of $\gamma$-ray observations of extragalactic sources (see, e.g., Ref. [7] for a review). Very-high energy (VHE, $E>100$ GeV) $\gamma$-ray photons can be absorbed during their propagation by the extragalactic background light (EBL) via the $\gamma$-$\gamma$ pair production process ($\gamma$$\gamma\rightarrow e^{+}+e^{-}$) [41, 27]. The created pairs inverse Compton (IC) scatter photons from the cosmic microwave background (CMB) up to $\gamma$-ray energies, initiating an electromagnetic cascade.
If an IGMF is present, the pairs are deflected, which induces several observational signatures of the cascade emission, such as an additional spectral component and a $\gamma$-ray halo around otherwise pointlike sources. The absence of these features in $\gamma$-ray observations of active galactic nuclei (AGNs) with their jets closely aligned to the line of sight (so-called blazars) has led to lower limits on the IGMF in the range of 10-15–10-17 G e.g., [40, 50, 19, 49, 4, 5]. The results depend on various assumptions such as the intrinsic VHE spectral shape, the AGN duty cycle, as well as the coherence length of the IGMF. Recently, a lower bound of the order of $\sim$ 10-17 G on IGMF strength was derived, which took into account the variability pattern of the source in the VHE band [3].
The discovery of the VHE emission from gamma-ray bursts (GRBs) [34, 1] offers a unique opportunity to constrain the IGMF. A non-zero IGMF would cause a time delay of the cascade emissions, producing observable $\gamma$-ray emission even after the GRB afterglow has ceased [43, 45, 37, 38, 48]. This approach has the advantage that it does not require assumptions on the source’s duty cycle or variability pattern as these quantities are directly determined for GRBs.
Several attempts have been made to measure the IGMF using this pair-echo emission using the first GRB detected in the VHE band, GRB 190114C [55, 23, 17, 54]. The authors of Ref. [55] excluded IGMF strengths lower than $\sim$ $10^{-19.5}$ G for a coherence length $\leq$ 1 Mpc, which was however, not confirmed in other studies [23, 17]. Instead, no limits could be derived. These discrepancies can be explained with the different assumptions made for the intrinsic VHE spectral shape of the GRB and shifting from semi-analytical to cosmic-ray propagation codes for the calculation of the cascade spectral energy distributions (SEDs). Therefore no IGMF constraint can be inferred from this GRB.
@@ II The VHE intrinsic spectrum of GRB 221009A
The GRB 221009A was an extremely energetic ($E_{\gamma}^{\rm iso}\sim 10^{55}$ erg) event detected with the Fermi gamma-ray burst monitor (GBM) on 2022 October 9 at $T_{0}=13:16:59.99$ UT [52]. With a redshift $z=0.151$, corresponding to a luminosity distance $D_{L}\sim 720$ Mpc, its isotropic luminosity is estimated to be $L_{\gamma}^{\rm iso}\simeq 10^{53}$ erg s-1 [18]. The event has been subsequently detected with the Fermi LAT in the GeV domain [11], and by the Swift-BAT and XRT instruments in X-rays [20]. Observations started 53 and 55 min after $T_{0}$, respectively.
At the trigger time $T_{0}$, the event was in the field of view of LHAASO, which detected more than $6.4\times 10^{4}$ photons within the first 3000 seconds at $E_{\gamma}>200$ GeV [32]. LHAASO is a hybrid array of VHE detectors, composed of both water Cherenkov photo-multipliers (WCDA) and solid-state scintillators (KM2A) that are sensitive to $\gamma$ rays above a few TeV [14, 33]. The intrinsic differential spectra in the range 200 GeV-7 TeV, derived within five temporal bins ranging from $T_{0}+5$ s to $T_{0}+1774$ s, do not show any evidence of energy cutoff up to 5 TeV [32]. The spectra are compatible with power laws (PLs) with indexes in the range 2.1–2.5. The KM2A observations reveal $\gamma$ rays with energies up to 13 TeV [13].
@@
Figure 2: Left: simulated SEDs of the pair-echo emission expected between 3.3 and 365 days after $T_{0}$ with the same injection spectrum as in Fig. 1 compared with the Fermi-LAT upper limits. Right: likelihood profiles sampling the magnetic field for different time intervals. The astrophysical GRB afterglow model assumes an ad hoc power law spectrum with $\Gamma=2$. Solid lines represent cubic spline interpolations between the derived likelihood values. The results were derived using the same set of simulations as in Fig. 1.
@@ III Simulation of the pair-echo emission
To produce the pair-echo SEDs in different time windows we employ the Monte-Carlo code CRPropa 3 (version 3.2), which simulates the development of the pair-echo emission given a primary VHE photon spectrum injected in the intergalactic medium [6]. The source is located at the center of a sphere with radius $D$ equal to the comoving distance to the source [^1] A redshift $z=0.151$ is sufficiently nearby for a reliable EBL implementation within CRPropa3, see Figure B.1 in [30].. We assume that the source emits VHE photons in the range 0.1-30 TeV according to Eq. (1), and re-weight the SED accordingly when a cutoff is adopted. These photons are injected considering an aperture cone of 1.6, as inferred from the break in the LHAASO light curve [32]. The choice of the cone aperture is based on the GRB emission model adopted by the LHAASO collaboration. We note that a larger opening angle could affect the cascade flux only for large time delays. However, for the time delays considered in this work, this parameter does not significantly impact the results.
Examples of the derived pair-echo light curves and SEDs are shown in Figs. 1 and 2. In Fig. 1, we show the cascade flux for $E>100$ MeV as a function of time for different IGMF strengths for a cutoff in the intrinsic spectrum of 7 TeV. The left panel of Fig. 2 shows the SED averaged between 3.3 days and 365 days. The lower value of 3.3 days corresponds to the last detection from the GRB as reported by Fermi LAT [9]. As expected, both figures show that the cascade flux drops with increasing magnetic field, since the time delay of the cascade photons increases with increasing IGMF strength. Furthermore, the peak of the cascade flux moves towards higher energies with increasing IGMF. This is caused by the energy dependence of the deflection angle of the pairs: for the same magnetic field $B$, the deflection angle is proportional to $E^{-2}$ for large correlation lengths [39].
This is due to the longer time delay of low energy photons as low energy pairs have a larger Larmor radius for higher IGMF values. We provide further details on the CRPropa simulations in Appendix A.
IV Fermi-LAT limits on the pair echo emission
The Fermi LAT is a pair-conversion telescope that detects $\gamma$ rays from 20 MeV to beyond 300 GeV [8]. We select events between 100 MeV and 1 TeV, within a $10^{\circ}\times 10^{\circ}$ region centered on GRB 221009A, with a zenith angle cut at $90^{\circ}$. We divide the dataset in several time bins following the main analysis of the Fermi-LAT Collaboration [9], expanding it up to one year after $T_{0}$ (Appendix B).
The likelihood of observing the data $\mathcal{D}_{ij}$ in the $i$th time bin and $j$th energy bin follows a Poisson likelihood $\mathcal{L}(B,\bm{\theta}_{i}|\mathcal{D}_{ij})$. This assumes an IGMF strength $B$ and spectral parameters $\bm{\theta}$ for both the background sources and potential GRB afterglow emission. Background sources can be nearby point sources or diffuse backgrounds. The log-likelihood ratio test can then be written as
@@ IV Fermi-LAT limits on the pair echo emission
where $\hat{\hat{\bm{\theta}}}_{i}$ denotes the spectral parameters maximizing $\mathcal{L}_{i}$ for a fixed IGMF field strength in the $i$th time bin, whereas $\hat{\bm{\theta}}_{i}$ and $\hat{B}$ are the parameters that maximize $\mathcal{L}_{i}$ unconditionally. Note that, for time bins earlier than $T_{0}+3.3\,$ days, we additionally included a point source with a power law spectrum with index $\Gamma=2$ and free normalization to account for the GRB afterglow emission (see Appendixes B and C).
We show the results of the log-likelihood ratio test in the right panel of Fig. 2. No IMGF strength is preferred. Instead, we are able to rule out IGMF strengths at 95 % confidence for which $\lambda>2.71$. Combining all time and energy bins, this excludes $B<2.5\times 10^{-17}$ G. The most stringent constraint comes from the latest time bin (up to $T_{0}+1$ yr), which rules out $B<2.2\times 10^{-17}$ G. Constraints from the individual time bins are reported in Appendix B.
@@ V Discussion and conclusions
Our results depend on the chosen high-energy cutoff of the injection spectrum of VHE $\gamma$ rays as well as the assumed GRB afterglow emission in the early time bins. Increasing the cutoff energy to 13 TeV, corresponding to the highest energy reported with the LHAASO KM2A detectors rules out $B<3.5\times 10^{-17}$ G. This is expected, as photons injected at even higher energies will increase the amount of pair-echo photons in the Fermi-LAT energy band. If no cutoff is applied to the log-parabola spectrum the constraints further improve and rule out an IGMF with $B<5\times 10^{-17}$ G. These results assume that an afterglow component can account for the GeV signal, fitting a power law with $\Gamma=2$ to the flux observed by the LAT in each time bin. To estimate the impact of assuming such an ad hoc component, we alternatively test a scenario where we replace the simple power law model with a physically motivated SSC model. This model is set to reproduce the LHAASO data, and only employed to predict the GeV flux in the $i$th time bin (Appendix C). The limits weaken marginally, and an IGMF with $B<1.4\times 10^{-17}\,$G is ruled out for a high-energy cutoff at 7 TeV. Finally, we also considered the scenario of no astrophysical afterglow. In this case, a preference for the cascade is found in the early time bins in order to explain the Fermi LAT detection. However, different time bins prefer different values of the IGMF, which are all incompatible with the constraints from the latest time bin extending up to $T_{0}+1\,$yr. Therefore, we do not regard this as evidence for an IGMF.
Figure 3: The IGMF parameter space ($B$, $\ell_{B}$). Theoretically preferred regions by electroweak (EWPT) or QCD phase transitions (QCDPT), or by galactic winds and induced by cosmic-ray (CR) streaming are shown in blue [22]. Previous constraints using LAT observations of GRB 221009A, as well as constraints from $\gamma$-ray observations of blazars are shown as gray lines and gray shaded regions with arrows. Our constraints at the 95 % confidence level are shown in orange.
In principle, the choice of the EBL model could affect the amount of cascade flux. In this work, we used the EBL model by [26]. However, as we show in Appendix A, the cascade flux level does not vary significantly across different EBL models, so the IGMF limit is not sensitive to the particular choice of model.
@@
Our results are significantly more constraining (by more than an order of magnitude) than previous analyses also using Fermi-LAT observations of GRB 221009A [29, 24, 53], see Fig. 3. The most significant factor concerns the time window considered for searching the pair echo. Previously, the cascade signal was searched for up to 20 days [53], 90 days [24] or eight months [29] after $T_{0}$. Furthermore, in our analysis we make full use of both spectral and temporal information given the chosen time binning.
Given the exceptionally high flux in the VHE band, it is plausible that Fermi-LAT observations could still be sensitive to the pair-echo emission at later times of more than one year. We assess this possibility by extending our simulations and extrapolating the LAT exposure up to five years. Given the energy-dependent dilution in time of the signal, we find that the average flux expected within the 3–15 GeV band will tighten the constraints if several years of exposure are added, despite not reaching an additional order of magnitude (see Appendix B).
The lower bound on the IGMF strength derived in this work holds as long as the correlation length $\ell_{B}$ of the IGMF is larger than the cooling distance of the electron-positron pairs for IC scattering, $l_{IC}\simeq 0.7\left(E_{\gamma}/~\rm{TeV}\right)^{-1}$ for primary $\gamma$ rays with energy $E_{\gamma}$. Considering $E_{\gamma}\sim 10\,$TeV, we find $\ell_{B}<l_{\mathrm{IC}}$ for $\ell_{B}\sim 0.1\,$Mpc, under the assumption that the electron and positron have both half the energy of the $\gamma$ ray. For lower values of the coherence length, the lower bound on the IGMF strength should scale as $\ell_{B}^{-1/2}$ for a Kolmogorov turbulence spectrum [39], as electron-positron stochastic deflections can be described as diffusion in angle (see Ref. [15] for the scaling with different turbulence spectra). Our result represents the best constraint so far on the IGMF employing the pair-echo technique, comparable to those obtained by studying blazars [3]. This opens the door for IGMF studies by means of TeV-detected GRBs to be as competitive as those done on persistent extragalactic VHE sources, while not suffering from assumptions on the duty cycles [5]. Moreover, our results are not affected by the possible development of plasma instabilities in the electron-positron beam in voids [12]. It has been argued that such instabilities could lead to an energy loss of the pairs on timescales shorter than the IC cooling time. However, the time required for such instabilities to grow is 1 order of magnitude larger than the time activity of the GRB 221009A, even with its large TeV luminosity [12, 32].